OpenStudy (anonymous):

Elijah forgot his pass code for a certain website. He knows it is made up of 4 digits and no digit can be repeated. If he randomly chooses 4 different digits, what is the probability that he chooses his pass code?

OpenStudy (kropot72):

The number of permutations of the digits 0 to 9 taken 4 at a time without repetitions is \[10P4=\frac{10!}{(10-4)!}=10\times 9\times 8\times 7\] When you have found this number of permutations take the reciprocal to find the required probability.